# Point-JEPA: A Joint Embedding Predictive Architecture for Self-Supervised Learning on Point Cloud

### v3

@@ Abstract
Recent advancements in self-supervised learning in the point cloud domain have demonstrated significant potential. However, these methods often suffer from drawbacks, including lengthy pre-training time, the necessity of reconstruction in the input space, or the necessity of additional modalities. In order to address these issues, we introduce Point-JEPA, a joint embedding predictive architecture designed specifically for point cloud data. To this end, we introduce a sequencer that orders point cloud tokens to efficiently compute and utilize tokens’ proximity based on their indices during target and context selection. The sequencer also allows shared computations of the tokens’ proximity between context and target selection, further improving the efficiency. Experimentally, our method achieves competitive results with state-of-the-art methods while avoiding the reconstruction in the input space or additional modality. Our code is available at https://github.com/Ayumu-J-S/Point-JEPA
@@ 1 Introduction
Sef-supervised learning is a paradigm that allows the model to learn a meaningful representation from unlabeled data. This allows the model to utilize a vast amount of unlabeled data and learn a strong representation, leading to advancements in natural language processing [9, 4, 20, 21] and 2D computer vision [2, 8, 12, 5, 11].
Only recently, we have seen the successful applications of self-supervised learning in the point cloud domain [17, 29, 27, 28, 7], achieving state-of-the-art results on downstream tasks. However, our initial investigation found that they require a significant amount of pre-training time as shown in Figure 1. This slow pre-training process can pose constraints in scaling to a larger dataset or deeper and more complex models, hindering the key advantage of sel-supervised learning; its capacity to learn a strong representation from a vast amount of data.
The recently proposed Joint-Embedding Predictive Architecture [14] and successful implementations of the architecture for pre-training a model [2, 3] show JEPA’s ability to learn strong off-the-shelf semantic representations. The idea behind JEPA is to learn a representation by predicting the embedding of the input signal, called target, from another compatible input signal, called context, with the help of a predictor network. This allows learning in the representation space instead of the input space, leading to efficient learning.
@@
Inspired by I-JEPA [2], we aim to apply Joint-Embedding Predictive Architecture in the point cloud domain. While this application introduces a promising direction for self-supervised learning in the point cloud domain, it also presents a unique challenge. Unlike an image, a point cloud is a set of points with no specific order. Regardless of the permutations applied to the set, the set presents represents the same object. This implies that the model consuming $N$ points needs to be invariant to $N!$ permutations of the order of the input data [18]. This unordered nature of the point cloud data makes the context and target selection of the data difficult especially if we aim to select spatially contiguous patches similar to I-JEPA [2]. In order to overcome this challenge while utilizing the full potential of Joint-Embedding Predictive Architecture for efficiency, we introduce Point-JEPA. To this end, we introduce a sequencer that orders the input sequence such that elements keep the spatial proximity when they are adjacent in the data. This allows the shared computation for proximity between the target and context, allowing efficiency while achieving a state of the performance with linear evaluation on ModelNet40 [26] as shown in Figure 1. Our method also achieves promising results in other downstream tasks showing the strong and transferable learned representation.
@@ 2 Point-JEPA Architecture
Our aim is to bring Joint-Embedding Predictive Architecture to the point cloud domain while measuring the efficient implementation. The overall framework, as shown in Figure 2, first converts the point cloud to a set of tokens, then the sequencer orders the tokens based on the spatial proximity, and Joint-Embedding Predictive Architecture is applied to the ordered tokens. We utilize a small PointNet [18] architecture for encoding the grouped points and standard Transformer [25] architecture for the context and target encoder as well as the predictor. It is important to note that our JEPA architecture operates on the token instead of patches in order to share the point encoder network between context and target encoder for efficiency similar to Point2Vec [28].
Point Cloud Tokenizer.
@@
Similar to previous studies utilizing standard Transformer architecture on point cloud objects [17, 27, 7, 28], we adopt a point cloud tokenizer that embeds groups of points. Given a point cloud object, $c$ center points are first sampled using the farthest point sampling [19]. Then using the $k$-nearest neighbors algorithm, we sample $k$ closest points to each of $c$ center points. These point patches are then normalized by subtracting the center point coordinate from the coordinates of the points in the patches. This allows the separation between local structural information and the positional information of the patches. In order to embed the local point patches, we utilize a small PointNet [18] architecture. This ensures that the patch embedding, or token, remains invariant to any permutations of data feeding order of points within the patch. Specifically, this PointNet contains two sets of a shared MLP and a max-pooling layer. First, a shared MLP maps each point into a feature vector. Then, we apply max-pooling to these vectors and concatenate the result back to the original feature vector. Another shared MLP then processes these concatenated vectors and a max-pooling layer is applied to produce tokens.
Sequencer.
@@
Because of the previously observed advantages of spatial proximity in patches for targets and context, referred to as a block in I-JEPA [2], we aim to sample tokens that are spatially close to each other. As previously mentioned, point cloud data is permutation invariant to data feeding order, which implies that even if the indices of tokens are sequential, they might not be spatially adjacent. Moreover, our method involves selecting spatially contiguous M blocks of encoded embedding vectors as the target while ensuring that the context does not include the tokens corresponding to these embedding vectors (details in the next paragraph). In order to overcome these challenges, we introduce a sequencer that is applied after the tokenization of the points. This sequencer orders tokens based on their associated center points. The process begins with a chosen center point and its associated token. In each subsequent step, the center point closest to the one previously chosen and its associated token are selected. This is iterated until the sequencer visits all of the center points. The resulting arrangement of tokens is in a sequence where contiguous elements are also spatially contiguous in most cases. This allows the shared computation of spatial proximity between context and target selection. At the same time, this also allows simpler implementation for context and target selection. It is worth noting, however, that selecting two adjacent token indices does not always guarantee spatial proximity; there might be a gap between them. While this is true, the experiment results show that this iterative ordering is effective enough in our JEPA architecture.
Figure 2: Point-JEPA. The tokenization of point cloud patches is shown on the left and the joint embedding architecture is shown on the right.
Context and Target.
@@
Targets in Point-JEPA can be considered patch-level representations of the point cloud object, which the predictor aims to predict. As illustrated in Figure 2, the target encoder initially encodes the tokens conventionally, and we randomly select M possibly overlapping target blocks, which are sets of adjacent embedding vectors. We denote these sets as target blocks. Context, on the other hand, is the representation of the point cloud object which is passed to the predictor to facilitate the reconstruction of target blocks. In order to avoid trivial learning, we remove the corresponding tokens of the target blocks during the context selection step. Out of this subset of tokens, we select a set of tokens that are adjacent in their indices and denote it as a context block. Because of this, context may be formed from multiple sets of spatially contiguous patches while targets often consist of one “block” of contiguous patches as shown in Figure 3.
Predictor and Loss
The task of predictor given targets and context is analogous to the task of supervised prediction. Given a context as input along with a certain condition, it aims to predict the target representations. Here, the condition involves the mask tokens, which are created from shared learned parameters, as well as positional encoding created from center points associated with the targets. Because the predictor’s task is to predict the representation produced by the target encoder, the loss can be defined to minimize the disagreement between the predictions and targets. To this end, we utilize Smooth L1 loss to measure the dissimilarity between target embedding and predicted embedding. The parameters of the target encoder, as well as the point tokenizer, are updated via gradient-based optimization, while the parameters of the target encoder are updated via the exponential moving average of the context encoder.
@@ 3 Experiments
In this section, we first describe the details of self-supervised pre-training. Following that, we compare the performance of the learned representation to some of the top-performing self-supervised learning methods in the point cloud domain, especially those that utilize the ShapeNet [6] dataset in pre-training. We specifically evaluate the learned representation using linear probing, end-to-end fine-tuning, and a few-shot learning setting.
@@ 3.1 Self-Supervised Pre-training
We pre-train our model on training set of ShapeNet [6] following the previous studies utilizing the standard Transformer [25] architecture such as Point-MAE [17], Point-M2AE [29], PointGPT [7], and Point2Vec [28]. The dataset consists of 41952 3D point cloud instances created from synthetic 3D meshes from 55 categories. As previously mentioned, standard Transformer [25] architecture is used for the context and target encoder as well as the predictor. During pre-training, we set the number of center points to 64 and the group size to 32. The point tokenization is applied to the input point cloud containing 1024 points per object. We set the depth of the Transformer in the context and target encoder to 12 with the embedding width of 384 and 6 heads. For the predictor, we use the narrower dimension of 192 following I-JEPA [2]. The depth of the predictor is set to 6, and the number of heads is set to 6.

### v4

@@ Abstract
Recent advancements in self-supervised learning in the point cloud domain have demonstrated significant potential. However, these methods often suffer from drawbacks, including lengthy pre-training time, the necessity of reconstruction in the input space, or the necessity of additional modalities. In order to address these issues, we introduce Point-JEPA, a joint embedding predictive architecture designed specifically for point cloud data. To this end, we introduce a sequencer that orders point cloud tokens to efficiently compute and utilize tokens’ proximity based on their indices during target and context selection. The sequencer also allows shared computations of the tokens’ proximity between context and target selection, further improving the efficiency. Experimentally, our method demonstrates state-of-the-art performance while avoiding the reconstruction in the input space or additional modality. In particular, Point-JEPA attains classification accuracy of $\bm{93.7}\scriptstyle\pm 0.2$ for linear SVM on ModelNet40 and $\bm{92.9}\scriptstyle\pm 0.4$% on the ScanObjectNN OBJ-BG dataset, surpassing all other self-supervised models. Moreover, Point-JEPA also establishes new state-of-the-art performance levels across all four few-shot learning evaluation frameworks. Code is available at https://github.com/Ayumu-J-S/Point-JEPA
@@ 1 Introduction
The growing accessibility of affordable consumer-grade 3D sensors has led to the widespread adoption of point clouds as a preferred data representation for capturing real-world environments. However, the existing point cloud understanding approaches [15] mostly rely on supervised training which requires time-consuming and labor-intensive manual annotations to semantically understand 3D environments. On the other hand, self-supervised learning (SSL) is an evolving paradigm that allows the model to learn a meaningful representation from unlabeled data. The success of self-supervised learning in advancing natural language processing and 2D computer vision has motivated its utilization in the point cloud domain for achieving state-of-the-art results on downstream tasks [19]. However, our initial investigation found that they require a significant amount of pre-training time as shown in Figure 1. The slow pre-training process can pose constraints in scaling to a larger dataset or complex and deeper models, hindering the key advantage of self-supervised learning; its capacity to learn a strong representation from a vast amount of data.
The successful implementations of Joint-Embedding Predictive Architecture (JEPA) [20] for pre-training a model [2, 4] show JEPA’s ability to learn strong semantic representations without the need for fine-tuning. The idea behind JEPA is to learn a representation by predicting the embedding of the input signal, called target, from another compatible input signal, called context, with the help of a predictor network. This allows learning in the representation space instead of the input space, leading to efficient learning. Inspired by I-JEPA [2], we aim to apply Joint-Embedding Predictive Architecture in the point cloud domain which introduces a promising direction for self-supervised learning in the point cloud understanding. However, unlike images that represent the same shape regardless of the permutations applied to it, unordered point clouds pose a unique challenge to the utilization of JEPA which requires permutation invariance [28]. The unordered nature of the point cloud data makes the context and target selection of the data difficult and inefficient, especially if we aim to select spatially contiguous patches similar to I-JEPA [2]. Therefore, we introduce Point-JEPA to overcome this challenge, while utilizing the full potential of Joint-Embedding Predictive Architecture for computational efficiency. Point-JEPA utilizes an efficient greedy tokenizer to assist the model to select point cloud tokens that are spatially adjacent. Empirical study indicates that Point-JEPA efficiently learns the semantic representations from point cloud data with fewer samples and faster pre-training. The specific contributions of this work are as follows.
  • We present a Joint-Embedding Predictive Architecture, called Point-JEPA, for point cloud SSL. Point-JEPA efficiently learns a strong representation from point cloud data without reconstruction in the input space or additional modality.
  • We propose a novel point cloud token ordering method for Joint-Embedding Predictive Architecture, utilizing a greedy algorithm based on spatial proximity.
2 Related Work
Recent advancements in self-supervised learning in 2D computer vision [2, 9, 6, 16, 14, 19, 37, 25] and natural language processing [10, 5, 30, 31] have inspired its application to point cloud processing. In this section, we review existing self-supervised learning works for point cloud learning.
Generative Learning
Generative learning models learn the representations by reconstructing the input data. These methods encourage the model to predict missing or transformed parts of the input, thereby gaining a deep understanding of the input’s intrinsic features. For example, based on a popular NLP model Bert [10], Point-Bert [40] introduces generative pretraining to the point cloud using a discrete variational autoencoder to transform the point cloud into discrete point tokens. However, this model heavily relies on data augmentation and suffers from the early leakage of location information, which makes pre-training steps relatively complicated and computationally expensive. To overcome this issue, Point-MAE [27] presents a lightweight, flexible, and computationally efficient solution by bypassing the tokenization and reconstructing the masked point cloud patches. On the other hand, PointGPT [8] introduces an auto-regressive learning model in the point cloud domain. A diffusion-based pre-training framework PointDiff [46] performs iterative denoising on the point cloud to extract hierarchical geometric prior. Alternatively, CluRender [23] utilizes differentiable rendering to learn transferable point-level features on a 3D point cloud. Such generative pre-training in the point cloud domain learns a robust representation; however, it suffers from computational inefficiency due to the reconstruction of the data in the input space.
Joint Embedding Architecture
Joint Embedding Architectures map the input data into a shared latent space that contains similar embeddings for semantically similar instances. These networks utilize regularization strategies such as contrastive learning and self-distillation to learn meaningful representations. Contrastive learning generates embeddings that are close for positive pairs and distant for negative pairs. For example, Du et al. [11] introduces a contrastive learning approach that treats different parts of the same object as negative and positive examples. Few studies [17, 38, 44] have applied contrastive learning to scene understanding from large-scale point clouds. For example, a recent work [38] strategically swaps objects between scenes and learns features robust to context variations.
Figure 2: Schematic renderings highlighting point cloud tokenization (Top left), point encoder (bottom left) and Point-JEPA (right). Point cloud patches are generated using furthest point sampling (FPS) [12] and $k$-nearest neighbor (KNN) methods, a mini PointNet (Point Encoder) is used to generate representative tokens which are subsequently fed to the JEPA architecture. We use standard Transformer [36] architecture for context ($f_{\theta}$) and target ($f_{\overline{\theta}}$) encoders as well as predictor ($p_{\phi}$).
Unlike contrastive learning, a self-distillation network employs two identical networks with distinct parameters, commonly known as the teacher and student, where the teacher guides the student by providing its predictions as targets. For example, in DCGLR [13], the teacher receives a global set obtained by cropping the original point cloud and the student receives the global set along with a local set obtained by cropping the point cloud with a larger cropping ratio. Similarly, inspired by [3], in Point2Vec [41], the teacher receives the patches of point clouds while the student receives a subset of these patches. Further, a shallow Transformer learns meaningful and robust representation from the masked positional information and the contextualized embedding from the partial-view input. In self-distillation networks, no reconstruction in the input space results in faster training than in generative models. However, as shown in Fig. 1, it requires longer training to learn a meaningful representation. On the other hand, contrastive learning excels in performance; however, its effectiveness highly depends on the careful selection of positive and negative samples as well as the data augmentation techniques to ensure transferable representations for downstream tasks [19].
Joint Embedding Predictive Architecture (JEPA)
A self-supervised learning architecture JEPA [26] learns representation using a predictor network that predicts one set of encoded signal $y$ based on another set of encoded signal $x$, along with a conditional variable $z$ that controls the prediction. In the predictor network, encoders initially process both the target and the context signals to represent them in embedding space. Conceptually JEPA has a large similarity to generative models which are designed to reconstruct masked part of the input. However, instead of directly operating on the input space, JEPA makes predictions in the embedding space. This allows the elimination of unnecessary input details to focus on learning meaningful representations. As a result, the model can abstract and represent the data more efficiently.
Closely related to our work, the specific application of the architecture in the image domain can be seen in I-JEPA [2]. In this work, the context signal is created by encoding a block of patches while the target signals are created from sampling blocks from encoded embedding vectors. Experiments show faster convergence of I-JEPA to learn highly semantic representation. Therefore, to ensure faster pretraining in self-supervised learning for point cloud understanding, we aim to apply JEPA on point cloud data.
@@ 3 Point-JEPA Architecture
In this section, we describe our JEPA architecture for pretraining in the point cloud domain. Our goal is to adapt JEPA [20] for use with point cloud data while evaluating its performance and implementation efficiency. The overall framework, as shown in Fig. 2, first converts the point cloud to a set of tokens, then a greedy sequencer arranges the tokens in sequence based on their spatial proximity to each other, and Joint-Embedding Predictive Architecture is applied to the ordered tokens. We utilize a mini PointNet [28] architecture for encoding the grouped points and standard Transformer [36] architecture for the context and target encoder as well as the predictor. It is important to note that our JEPA architecture operates on the tokens instead of patches in order to share the point encoder network between context and target encoder for efficiency similar to Point2Vec [41].
3.1 Point Cloud Tokenizer
@@
Similar to previous studies utilizing standard Transformer architecture on point cloud objects [27, 40, 8, 41], we adopt a point cloud tokenizer that embeds groups of points. Given a point cloud object $P\subset\mathbb{R}^{3}$ consisting of $n$ points, $c$ center points are first sampled using the farthest point sampling [12]. Then we employ the $k$-nearest neighbors algorithm to identify and select the $k$ closest points surrounding each of the $c$ designated center points. These point patches are then normalized by subtracting the center point coordinate from the coordinates of the points in the patches. This allows the separation between local structural information and the positional information of the patches. In order to embed the local point patches to create tokens, we utilize a mini PointNet [28] architecture. This ensures that the patch embedding, or token, remains invariant to any permutations of data feeding order of points within the patch. Specifically, this PointNet contains two sets of a shared multi-layer perceptron (MLP) and a max-pooling layer as shown in Fig. 2. First, a shared MLP maps each point into a feature vector. Then, we apply max-pooling to these vectors and concatenate the result back to the original feature vector. Subsequently, a shared MLP processes these concatenated vectors, followed by a max-pooling operation to generate a set of representative tokens $T$ of $P$.
1
Input : Set of tokens, $T=\{t_{1},t_{2},..,t_{r}\}$
Output : Set of spatially contiguous tokens, $T^{\prime}$
2 Find the initial token $t=minCoordSum(T)$;
3 Set $T^{\prime}=\{t\}$;
4 $T=T\setminus\{t\}$;
5 Initialize $prev\_token=t$;
6 while $T\neq\emptyset$ do
7       Set $closest=\infty$ ;
8       for $t_{i}\in T$ do
9             $dis=\parallel center(prev\_token)-center(t_{i})\parallel$;
10             if $dis\leq closest$ then
11                   Set $closest=dis$;
12                   Set $index=i$;
13
14             end if
15
16       end for
17      $T^{\prime}=T^{\prime}\cup\{t\_index\}$;
18       $T=T\setminus\{t\_index\}$;
19       $prev\_token=t\_index$;
20
21 end while
22
Algorithm 1 Greedy token sequencer strategy
3.2 Greedy Token Sequencer
@@
Due to the previously observed benefits of having targets and context clustered together in close spatial proximity, a configuration known as a block in I-JEPA [2], we aim to sample tokens that are spatially close to each other. As previously mentioned, point cloud data is permutation invariant to data feeding order, which implies that even if the indices of tokens are sequential, they might not be spatially adjacent. Furthermore, our approach involves the selection of $M$ spatially contiguous blocks of encoded embedding vectors as the target while ensuring that the context does not include the tokens corresponding to these embedding vectors (details in the next paragraph). To address these challenges, we apply a greedy sequencer that is applied after the tokenization of the points similar to z-ordering applied in [8]. This sequencer orders tokens based on their associated center points (Algorithm 1). The process is initiated by selecting the center point with the lowest sum of coordinates ($minCoorSum(T)$) as the starting token, along with its associated embedding. In each subsequent step, the center point closest to the one previously chosen and its associated token are selected. This is iterated until the sequencer visits all of the center points. The resulting arrangement of tokens ($T^{\prime}=\{t_{1}^{{}^{\prime}},t_{2}^{{}^{\prime}},...,t_{r}^{{}^{\prime}}\}$) is in a sequence where contiguous elements are also spatially contiguous in most cases. This allows the shared computation of spatial proximity between context and target selection. At the same time, this also allows simpler implementation for context and target selection. It is worth noting, however, that selecting two adjacent token indices does not always guarantee spatial proximity; there might be a gap between them. While this is true, the experiment results show that this iterative ordering is effective enough in our JEPA architecture.
Additionally, this rather simple approach is parallelized across batches, making it more efficient for large datasets or point clouds. We can not only compute pairwise distances for all points within a batch in a single forward pass but also run the iterative process of selecting the next closest point across the batch simultaneously. This enables faster computation on modern GPUs, ensuring that the nearest points are selected efficiently while keeping the algorithm feasible even for large batch sizes.
3.3 JEPA Components
Context and Target
@@ 3.3 JEPA Components
Targets in Point-JEPA can be considered patch-level representations of the point cloud object, which the predictor aims to predict. As illustrated in Figure 2, the target encoder initially encodes the tokens conventionally, and we randomly select $M$ possibly overlapping target blocks, which are sets of adjacent embedding vectors. Specifically, we define $y(i)={\{y_{j}\}}_{j\in B_{i}}$ as the $i^{\text{th}}$ target representation block, where $B_{i}$ denotes the set of mask indices for the $i^{\text{th}}$ block. Here, we denote the encoded embedding vectors as $y=\{y_{1},y_{2},\ldots y_{n}\}$, where $y_{k}=f_{\overline{\theta}}(t_{k}^{{}^{\prime}})$ is the representation associated with the $k^{th}$ centre point. It is important to note that masking for the target is applied to the embedding vectors, which are derived from the tokens that have passed through the transformer encoder $f_{\overline{\theta}}$. This ensures a high semantic level for the target representations.
@@
Context, on the other hand, is the representation of the point cloud object which is passed to the predictor to facilitate the reconstruction of target blocks. Unlike target blocks, masking is applied to the point cloud tokens during the creation of context blocks. This allows the context-encoder $f_{\theta}$ to represent the uncertainties in the object’s representations when certain parts are masked. Specifically, we first select a subset of tokens $\hat{T}\subseteq T^{\prime}$ that are spatially contiguous. These selected tokens are then fed to the context-encoder $f_{\theta}$ to generate a context block $x=\{x_{j}\}_{j\in B_{x}}$. To prevent trivial learning, we also ensure that the indices of tokens chosen for the context differ from those for the targets. Furthermore, the tokens $T^{\prime}$ are sorted such that tokens adjacent in the data feeding order are also spatially close. This organization simplifies the selection of contiguous target and context blocks, despite the aforementioned complexities of point cloud data representation.
Predictor
The task of predictor $p_{\phi}$ given targets $y$ and context $x$ is analogous to the task of supervised prediction. Given a context as input $x$ along with a certain condition, it aims to predict the target representations $y$. Here, the condition involves the mask tokens, which are created from shared learned parameters, as well as positional encoding, created from centre points associated with the targets. That is
$\hat{y}(i)=\{\hat{y}_{j}\}_{j\in B_{i}}=p_{\phi}\left(x,\{m_{j}\}_{j\in B_{i}}\right)$ (1)
where $p_{\phi}\left(\cdot,\cdot\right)$ denotes the predictor and $\{m_{j}\}_{j\in B_{i}}$ denotes the mask token created from shared learnable parameter and positional encoding created from centre points.
Loss
Because the predictor’s task is to predict the representation produced by the target encoder, the loss can be defined to minimize the disagreement between the predictions and targets as follows.
$\frac{1}{M}\sum_{i=1}^{M}D(\hat{y}(i),y(i))=\frac{1}{M}\sum_{i=1}^{M}\sum_{j\in B_{i}}\mathcal{L}(\hat{y}_{j},y_{j})$ (2)
Similar to Point2Vec [41], we utilize Smooth L1 loss to measure the dissimilarity between each corresponding element of the target and predicted embedding because of its ability to be less sensitive to the outliers.
Parameter Update
We utilize AdamW [22] optimizer with cosine learning decay [21]. The target encoder and context encoder initially have identical parameters. The context encoder’s parameters are updated via backpropagation, while the target encoders’ parameters are updated using the exponential moving average of the context encoder parameters, that is $\overline{\theta}\xleftarrow{}\tau\overline{\theta}+(1-\tau)\theta$ where $\tau\in[0,1]$ denotes the decay rate.
@@ 4 Experiments
In this section, we first describe the details of self-supervised pre-training. Further, we compare the performance of the learned representation to the state-of-the-art self-supervised learning methods in the point cloud domain that utilizes the ShapeNet [7] dataset in pre-training. We specifically evaluate the learned representation using linear probing, end-to-end fine-tuning, and a few-shot learning setting. Finally, ablation experiments are conducted to gain insights into the principal characteristics of Point-JEPA.
@@ 4.1 Self-Supervised Pre-training
We pre-train our model on training set of ShapeNet [7] following the previous studies utilizing the standard Transformer [36] architecture such as Point-MAE [27], Point-M2AE [42], PointGPT [8], and Point2Vec [41]. The dataset consists of 41952 3D point cloud instances created from synthetic 3D meshes from 55 categories. The standard Transformer [36] architecture is used for the context and target encoder as well as the predictor. During pre-training, we set the number of center points to 64 and the group size to 32. The point tokenization is applied to the input point cloud containing 1024 points per object. We set the depth of the Transformer in the context and target encoder to 12 with the embedding width of 384 and 6 heads. For the predictor, we use the narrower dimension of 192 following I-JEPA [2]. The depth of the predictor is set to 6, and the number of heads is set to 6. Our experiments are conducted on NVIDIA RTX A5500 and NVIDIA A100 SXM4. We note that our method only takes 7.5 hours on RTX A5500 for pretraining (see Fig. 1) which is less than half of that of PointGPT [8] and about 60% of Point2Vec [41] time requirement for pretraining. Adhering to the standard convention, we use overall accuracy for classification tasks and mean IoU for part segmentation tasks.
Table 1: End-to-End Classification. Overall accuracy on ModelNet40 [39] and ScanObjNN [34] with end-to-end fine-tuning. We specifically compare our methods to the method utilizing standard Transformer architecture pre-trained on ShapeNet [7] with only point cloud (no additional modality).
Overall Accuracy
Method Reference ModelNet40 ScanObjNN
#Points +Voting -Voting #Points OBJ-BG OBJ-ONLY OBJ-T50-RS
Point-BERT [40] CVPR2022 1k 93.2 92.7 1k 87.4 88.1 83.1
Point-MAE [27] ECCV2022 1k 93.8 93.2 2k 90.0 88.3 85.2
Point-M2AE [42] NeurIPS2022 1k 94.0 93.4 2k 91.2 88.8 86.4
Point2Vec [41] GCPR2023 1k 94.8 94.7 2k 91.2 90.4 87.5
PointGPT-S [8] NeurIPS2023 1k 94.0 2k 91.6 90.0 86.9
PointDiff [46] CVPR2024 1k 93.2 91.9 87.6
Point-JEPA (Ours) - 1k $94.1\scriptstyle\pm 0.1$ $93.8\scriptstyle\pm 0.2$ 2k 92.9 ±$\scriptstyle\pm 0.4$ $90.1\pm 0.2$ $86.6\pm 0.3$